# Nonlocal stress-energy tensor in time-dependent gravitational backgrounds

### v1

@@ Abstract
We analyze the renormalized stress-energy tensor (RSET) of a massless quantum scalar field in time-dependent gravitational backgrounds. Starting from its formal expression obtained within the covariant perturbative expansion to lowest order in the curvature, we evaluate the RSET in an arbitrary number of dimensions in terms of coordinate-space distributions. For time-dependent spherically symmetric spacetimes, we derive a multipole expansion and determine its asymptotic behavior. We find that the RSET is locally nonvanishing at null infinity and depends on the detailed dynamics of the collapsing body. However, the total emitted energy vanishes at this order, meaning that the leading contribution does not account for the energy density of the created particles. Nevertheless, by enforcing stress-tensor conservation up to second order in the curvature, we show that the total radiated energy can be extracted from the first-order RSET. Finally, we compute the induced quantum corrections to the metric at large distances, which display several unexpected features.
@@ I Introduction
Quantum field theory in curved spacetime provides the theoretical framework to describe quantum effects of matter fields in the presence of a classical gravitational background. Although gravity itself is not quantized, the presence of quantum fields can lead to important physical phenomena, such as the Hawking radiation of black holes, the generation of cosmological perturbations during inflation, and the appearance of vacuum polarization effects in strong gravitational fields. These processes, first analyzed in the 1970’s, revealed that the vacuum of a quantum field depends on the global properties of the spacetime, establishing a deep connection between general relativity, thermodynamics, and quantum theory [1, 2, 3, 4, 5].
@@
A particularly powerful and conceptually clear approach is to perform a covariant perturbative expansion of the effective action in powers of the curvature [6, 7, 8, 9]. When one considers as quantum fields the fluctuations around a background metric, the resulting expression can be interpreted as an effective field theory of quantum gravity at low energies [10]. The quantum corrections to the action appear as higher-order curvature terms, which are in general nonlocal and are suppressed by powers of the Planck scale. This perspective, provides a systematic way to compute quantum corrections to classical gravitational phenomena without assuming a full theory of quantum gravity. Within this framework, the nonlocal part of the effective action captures the long-distance propagation of quantum fluctuations, leading to nontrivial effects such as $1/r^{3}$ corrections to Newton’s potential [11, 12, 13] and running of gravitational couplings [14, 15].
@@
Conventions. We use Planck units and adopt the mostly plus convention for the metric in the Lorentzian setup (,+,+,)(-$\dots$). We also define the Ricci tensor as $R_{\mu\nu}:=R^{\rho}{}_{\mu\rho\nu}$, the Ricci scalar $R=R^{\mu}{}_{\mu}$, and $g=\lvert\mathrm{det}\,g_{\mu\nu}\rvert$. As usual, Greek indices $(\mu,\nu,\dots)$ run over spacetime indices, while Latin indices $(i,j,\dots)$ correspond just to the spatial sector. For Fourier transforms we adopt the convention
@@ II RSET for a scalar quantum field in curved spacetime
where $\xi$ is the nonminimal coupling to the curvature. The calculation of the effective action $W$ associated with this quantum field is a complicated task. It can be carried out using a covariant perturbation theory [6, 7, 8], yielding a nonlocal expansion in powers of the curvature. Up to second order, and after performing a Wick rotation, the renormalized nonlocal part of the in–out effective action reads [6, 16]
@@
$\braket{0_{\text{in}}|T_{\mu\nu}|0_{\text{in}}}:=\braket{T_{\mu\nu}}=\left.-\frac{2}{\sqrt{g}}\,\frac{\delta W}{\delta g^{\mu\nu}}\right|_{\text{retarded prescription}}\,.$ (5)
@@ III Nonlocal operators as distributions in coordinate space
where the limit $\epsilon\to 0^{+}$ is taken at the end. Using the identity from Ref. [20],
@@
where $\Theta$ is the Heaviside step function, and [1]
@@
where the inner integral contributes only a numerical factor depending on $D$ and $\alpha$. Evaluating it analytically [21] yields
@@
To proceed, we use results from distribution theory [22]. When $\alpha-D/2$ is not an integer,
@@
When $\alpha-D/2=-n$ with $n\in\mathbb{N}$, one instead has [22]
@@
where $\delta^{(m)}$ is the $m$-th derivative of the Dirac delta, and Eq. (15) becomes
@@
The factor $\Theta(t-t^{\prime})\delta^{(D/2-\alpha-1)}(\sigma)$ restricts the support of the distribution to the surface of the past light cone, thus satisfying Huygens’ principle [23].
@@
which is supported on the past light cone in all dimensions. This result agrees with the odd-dimensional case obtained in Ref. [23] and reproduces the four-dimensional expression derived for static backgrounds in Ref. [24], as well as the time-dependent generalization in Ref. [25].
@@ IV.1 Leading order of the RSET in time-dependent scenarios
To summarize, we have found that the RSET at linear order in curvature powers carries an asymptotic energy flux, but the total energy in $\mathscr{I}^{+}$ vanishes upon integration over $u$ (this behavior was previously described in Ref. [26]). However, the information of particle creation is already present in the effective action, from which the RSET was derived [16]. In Section V we will clarify this somewhat puzzling situation, showing that the total energy change in the quantum field due to the time dependent gravitational field can indeed be extracted from the linear-order RSET.
@@ IV.2 The four dimensional case
This result highlights the inherently nonlocal character of the renormalized stress–energy tensor (RSET): a localized source of curvature gives rise to a nontrivial stress–energy distribution of the quantum field throughout spacetime. In contrast with the time-independent case [24, 25, 27, 28], the leading-order term in Eq. (41) depends explicitly on both the spatial profile and the temporal history of the curvature source, rather than solely on its total associated mass. An explicit example will be analyzed in Sec. VI. This dependence of the RSET—and consequently of the quantum-corrected metric—on the detailed spatial structure of the source is often referred to as quantum hair.
@@ IV.3 Quantum corrections to the metric tensor
In previous studies (see, for instance, Ref. [24]), it was shown that for time-independent backgrounds with compact sources in four dimensions, the RSET falls off as $M/r^{5}$, with $M$ being the mass characterizing the compact curvature source, leading to quantum corrections to the metric tensor with an asymptotic behavior $M/r^{3}$. In contrast, in Sec. IV.2 we have shown that, once time dependence is taken into account, the leading-order contribution to the RSET depends on the second time derivative of the curvature scalar and exhibits an asymptotic behavior proportional to ${M^{\prime\prime}_{0}}^{(R)}(t-r)/r^{2}$ (see Eq. (41)), where $M^{(R)}_{0}(t-r)$ is the function defined in Eq. (40). We now examine how this contribution gives rise to quantum corrections to the metric tensor.
In linearized gravity, writing $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$ in Cartesian coordinates, the Einstein equations read
$\Box\bar{h}_{\mu\nu}=-16\pi\left(T_{\mu\nu}^{\textrm{(cl)}}+\braket{T_{\mu\nu}}\right),\qquad\bar{h}_{\mu\nu}=h_{\mu\nu}-\frac{1}{2}h\,\eta_{\mu\nu},$ (42)
where $h=\eta^{\mu\nu}h_{\mu\nu}$, $\Box$ is the flat-space d’Alembertian, and the perturbation is assumed to satisfy the de Donder gauge condition, $\partial^{\mu}\bar{h}_{\mu\nu}=0$. Here, $T_{\mu\nu}^{\textrm{(cl)}}$ denotes the classical part of the stress-energy tensor, which acts as the curvature source of the background. To compute the back-reaction on $h_{\mu\nu}$, we use the semiclassical gravity approach and break $h_{\mu\nu}$ into two pieces:
@@
where $h_{\mu\nu}^{\mathrm{(cl)}}$ denotes the classical part, while $h_{\mu\nu}^{\mathrm{(q)}}$ denotes the quantum one. Since $h_{\mu\nu}^{\mathrm{(cl)}}$ is the solution of the classical part of Einstein equations, then Eq. (42) implies that $h_{\mu\nu}^{\mathrm{(q)}}$ satisfies
$\Box\bar{h}_{\mu\nu}^{\mathrm{(q)}}=-16\pi\braket{T_{\mu\nu}}.$ (44)
Solving Eq. (44) requires knowledge of the RSET throughout spacetime, due to the nonlocal nature of the Green’s function of the d’Alembert operator. To capture the asymptotic local contribution of $\braket{T_{\mu\nu}}$ to $\bar{h}_{\mu\nu}^{\mathrm{(q)}}$, we adopt the ansatz
$\bar{h}^{t}{}_{t}^{\mathrm{(q)}}=\bar{h}^{t}{}_{r}^{\mathrm{(q)}}=\bar{h}^{r}{}_{r}^{\mathrm{(q)}}=\frac{1}{r}\log\left(\mu r\right)q(t-r),$ (45)
where $\mu$ is an arbitrary parameter. The function q(tr)q(t-r) in Eq. (45) is then determined by substituting the ansatz into Eq. (44). Note that the proposed $\bar{h}_{\mu\nu}^{\mathrm{(q)}}$ in is traceless, which implies $\bar{h}_{\mu\nu}^{\mathrm{(q)}}=h_{\mu\nu}^{\mathrm{(q)}}$. Taking the D’Alembertian of the ansatz in Eq. (45) we obtain
$\Box h^{\mu}{}_{\nu}^{\mathrm{(q)}}=\left(\begin{array}[]{cccc}\frac{2}{r^{2}}\dot{q}+\frac{1}{r^{3}}q&-\frac{2}{r^{2}}\dot{q}-\frac{2\log\left(\mu r\right)+1}{r^{3}}q&0&0\\ \frac{2}{r^{2}}\dot{q}+\frac{2\log\left(\mu r\right)+1}{r^{3}}q&-\frac{2}{r^{2}}\dot{q}-\frac{4\log\left(\mu r\right)+1}{r^{3}}q&0&0\\ 0&0&\frac{2\log\left(\mu r\right)}{r^{3}}q&0\\ 0&0&0&\frac{2\log\left(\mu r\right)}{r^{3}}q\end{array}\right),$ (46)
where the function $q$ and its derivative $\dot{q}$ are evaluated at trt-r. Direct comparison with Eq. (41) shows that leading-order terms $\sim 1/r^{2}$ match if
@@
while subleading terms should be fixed by adding subleading contributions to our ansatz. We have added the superscript $\mathrm{(cl)}$} to the curvature scalar $R$ in Eq. (47) to emphasize that it must be computed using the classical background metric. Observe that different choices of the parameter $\mu$ in Eq. (45) introduce vacuum-solution terms in the ansatz; therefore, it can be determined by imposing appropriate initial conditions.
The complete metric in the asymptotic region $r\to\infty$ outside a spherically-symmetric body of mass $M$ contains a classical part, which corresponds to the Schwarzschild metric expressed in isotropic coordinates (in order to satisfy the de Donder gauge condition), and the quantum correction. We obtain, in spherical coordinates,
$g^{\mu}{}_{\nu}=\delta^{\mu}{}_{\nu}+\underbrace{\frac{1}{r}\log\left(\mu r\right)q(t-r)\begin{pmatrix}-1&1&0&0\\ -1&1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}}_{=h^{\mu}{}_{\nu}^{\mathrm{(q)}}}+\underbrace{\frac{2M}{r}\begin{pmatrix}-1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix}}_{=h^{\mu}{}_{\nu}^{\mathrm{(cl)}}}+\mathcal{O}\left(\frac{\log(\mu r)}{r^{2}}\right),$ (48)
@@
At first glance, Eq. (48) might suggest a quantum-corrected gravitational wave. However, several crucial features distinguish it from its classical counterpart. First, the propagating mode in Eq. (48) appears even in spherically symmetric configurations, whereas classical gravitational waves are absent due to Birkhoff’s theorem. Second, the solution does not exhibit the usual transverse polarization. Third, its asymptotic behavior, $h^{\mu}{}_{\nu}^{\mathrm{(q)}}\sim\log(\mu r)/r$, differs from the standard $1/r$ decay of classical waves.
At this point, one might naively conclude that the energy transported by the quantum-corrected metric grows indefinitely at large $r$ because, schematically, the gravitational stress-energy pseudo-tensor $t_{\mu\nu}$ is proportional to $h_{\mu\nu}^{2}\sim\log^{2}(\mu r)/r^{2}$. However, following the definition given in Ref. [29],
@@
where the superscript $(1)$ denotes the corresponding linearized expression and the superscript $(2)$ the quadratic part in $h_{\mu\nu}=h_{\mu\nu}^{\mathrm{(cl)}}+h_{\mu\nu}^{\mathrm{(q)}}$, a straightforward calculation yields
$t^{\mu}{}_{\nu}=0+\mathcal{O}\left(\frac{\log(\mu r)}{r^{3}}\right)\,.$ (50)
The slowly decaying terms proportional to $\log^{2}(\mu r)/r^{2}$ are not present in the stress-energy pseudo-tensor associated with the corrected gravitational field. Hence, since $t^{tr}$ decays faster than $1/r^{2}$ as $r\to\infty$, no net energy flux is transported to infinity by the quantum-corrected metric.
@@
$\frac{\mathrm{d}^{2}S^{r}}{d\tau^{2}}$ $=\frac{2M}{r^{3}}\,$, (51)
$\frac{\mathrm{d}^{2}S^{\theta}}{d\tau^{2}}$ $=\frac{\mathrm{d}^{2}S^{\varphi}}{d\tau^{2}}=-\frac{M}{r^{3}}+\frac{1}{2r^{2}}\log\left(\mu r\right)\dot{q}(t-r)\,$,
where we have neglected the terms proportional to $q(t-r)\log(\mu r)/r^{3}$ for consistency with Eq. (46). The contributions proportional to the source mass $M$ are the well-known results corresponding to the Schwarzschild geometry, while the terms proportional to $\dot{q}$ in the angular components describe a tidal effect transverse to the propagation direction. Therefore, although the quantum correction to the metric has a polarization tensor that does not represent a genuine gravitational wave, its effect on massive test particles is analog to that of a classical wave.
Finally, we note that, at the formal level, the quantum-corrected metric dominates over its classical counterpart at large distances. However, this apparent dominance arises because we are working in Planck units. Restoring conventional units makes the quantum correction proportional to $\ell_{\text{Planck}}^{2}$, so that both the classical and quantum contributions become comparable only at unrealistically large distances. An explicit example illustrating this behavior is presented in Sec. VI.

### v2

@@ Abstract
We analyze the renormalized stress-energy tensor (RSET) of a massless quantum scalar field in time-dependent gravitational backgrounds. Starting from its formal expression obtained within the covariant perturbative expansion to lowest order in the curvature, we evaluate the RSET in an arbitrary number of dimensions in terms of coordinate-space distributions. For time-dependent spherically symmetric spacetimes, we derive a multipole expansion and determine its asymptotic behavior. We find that the RSET is locally nonvanishing at null infinity and depends on the detailed dynamics of the collapsing body. However, the total emitted energy vanishes at this order, meaning that the leading contribution does not account for the energy density of the created particles. Nevertheless, by enforcing stress-tensor conservation up to second order in the curvature, we show that the total radiated energy can be extracted from the first-order RSET. Finally, we compute the induced quantum corrections to the metric at large distances, which display several interesting features.
@@ I Introduction
Quantum field theory in curved spacetime provides the theoretical framework to describe quantum effects of matter fields in the presence of a classical gravitational background. Although gravity itself is not quantized, the presence of quantum fields can lead to important physical phenomena, such as the Hawking radiation of black holes, the generation of cosmological perturbations during inflation, and the appearance of vacuum polarization effects in strong gravitational fields. These processes, first analyzed in the 1970’s, revealed that the vacuum of a quantum field depends on the global properties of the spacetime, establishing a deep connection between general relativity, thermodynamics, and quantum theory [birrell1982, Fulling1989, Wald1995, Parker2009, Buchbinder2021].
@@
A particularly powerful and conceptually clear approach is to perform a covariant perturbative expansion of the effective action in powers of the curvature [avramidi1991, Barvinsky1987, Barvinsky1990, Barvinsky1994]. When one considers as quantum fields the fluctuations around a background metric, the resulting expression can be interpreted as an effective field theory of quantum gravity at low energies [Donoghue1994]. The quantum corrections to the action appear as higher-order curvature terms, which are in general nonlocal and are suppressed by powers of the Planck scale. This perspective, provides a systematic way to compute quantum corrections to classical gravitational phenomena without assuming a full theory of quantum gravity. Within this framework, the nonlocal part of the effective action captures the long-distance propagation of quantum fluctuations, leading to nontrivial effects such as $1/r^{3}$ corrections to Newton’s potential [Duff1973, Donoghue1993, Dalvit:1994] and running of gravitational couplings [Dalvit1994_2, Barvinsky2023].
@@
I.1 Conventions
We use Planck units and adopt the mostly plus convention for the metric in the Lorentzian setup (,+,+,)(-$\dots$). We also define the Ricci tensor as $R_{\mu\nu}:=R^{\rho}{}_{\mu\rho\nu}$, the Ricci scalar $R=R^{\mu}{}_{\mu}$, and $g=\lvert\mathrm{det}\,g_{\mu\nu}\rvert$. As usual, Greek indices $(\mu,\nu,\dots)$ run over spacetime indices, while Latin indices $(i,j,\dots)$ correspond just to the spatial sector. For Fourier transforms we adopt the convention
@@ II RSET for a scalar quantum field in curved spacetime
where $\xi$ is the nonminimal coupling to the curvature. The calculation of the effective action $W$ associated with this quantum field is a complicated task. It can be carried out using a covariant perturbation theory [avramidi1991, Barvinsky1987, Barvinsky1990], yielding a nonlocal expansion in powers of the curvature. Up to second order, and after performing a Wick rotation, the renormalized nonlocal part of the in–out effective action reads [avramidi1991, nosotros_cotton]
@@
$\braket{T_{\mu\nu}}:=\braket{0_{\text{in}}|T_{\mu\nu}|0_{\text{in}}}=\left.-\frac{2}{\sqrt{g}}\,\frac{\delta W}{\delta g^{\mu\nu}}\right|_{\text{retarded prescription}}\,.$ (5)
@@ III Nonlocal operators as distributions in coordinate space
where the limit $\epsilon\to 0^{+}$ is taken at the end. Using the identity from Ref. [lopez_nacir_fdm],
@@
where $\Theta$ is the Heaviside step function, and [birrell1982]
@@
where the inner integral contributes only a numerical factor depending on $D$ and $\alpha$. Evaluating it analytically [gradshteyn2007] yields
@@
To proceed, we use results from distribution theory [gelfand1964]. When $\alpha-D/2$ is not an integer,
@@
When $\alpha-D/2=-n$ with $n\in\mathbb{N}$, one instead has [gelfand1964]
@@
where $\delta^{(m)}$ is the $m$th derivative of the Dirac delta, and Eq. (15) becomes
@@
The factor $\Theta(t-t^{\prime})\delta^{(D/2-\alpha-1)}(\sigma)$ restricts the support of the distribution to the surface of the past light cone, thus satisfying Huygens’ principle [bollini-giambiagi].
@@
which is supported on the past light cone in all dimensions. This result agrees with the odd-dimensional case obtained in Ref. [bollini-giambiagi] and reproduces the four-dimensional expression derived for static backgrounds in Ref. [fdm_vacuum_polarization], as well as the time-dependent generalization in Ref. [calmet_grav].
@@ IV.1 Leading order of the RSET in time-dependent scenarios
To summarize, we have found that the RSET at linear order in curvature powers carries an asymptotic energy flux, but the total energy in $\mathscr{I}^{+}$ vanishes upon integration over $u$ (this behavior was previously described in Ref. [quantum_noise]). However, the information of particle creation is already present in the effective action, from which the RSET was derived [nosotros_cotton]. In Section V we will clarify this somewhat puzzling situation, showing that the total energy change in the quantum field due to the time dependent gravitational field can indeed be extracted from the linear-order RSET.
@@ IV.2 The four dimensional case
This result highlights the inherently nonlocal character of the renormalized stress–energy tensor (RSET): a localized source of curvature gives rise to a nontrivial stress–energy distribution of the quantum field throughout spacetime. In contrast with the time-independent case [fdm_vacuum_polarization, calmet_grav, universality_fdm, universality_anderson], the leading-order term in Eq. (41) depends explicitly on both the spatial profile and the temporal history of the curvature source, rather than solely on its total associated mass. An explicit example will be analyzed in Sec. VI. This dependence of the RSET—and consequently of the quantum-corrected metric—on the detailed spatial structure of the source is often referred to as quantum hair.
@@ IV.3 Quantum corrections to the metric tensor
In previous studies (see, for instance, Ref. [fdm_vacuum_polarization]), it was shown that for time-independent backgrounds with compact sources in four dimensions, the RSET falls off as $M/r^{5}$, with $M$ being the mass characterizing the compact curvature source, leading to quantum corrections to the metric tensor with an asymptotic behavior $M/r^{3}$. In contrast, in Sec. IV.2 we have shown that, once time dependence is taken into account, the leading-order contribution to the RSET depends on the second time derivative of the curvature scalar and exhibits an asymptotic behavior proportional to ${M^{\prime\prime}_{0}}^{(R)}(t-r)/r^{2}$ (see Eq. (41)), where $M^{(R)}_{0}(t-r)$ is the function defined in Eq. (40) We now examine how this contribution gives rise to quantum corrections to the metric tensor.
In linearized gravity, writing $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}$, the Einstein equations read
$G_{\mu\nu}^{(1)}=8\pi\left(T_{\mu\nu}^{\textrm{(cl)}}+\langle T_{\mu\nu}^{\textrm{(q)}}\rangle\right)\,$, (42)
where $G_{\mu\nu}^{(1)}$ denotes the linearized Einstein tensor and $T_{\mu\nu}^{\textrm{(cl)}}$ is the classical part of the stress-energy tensor, which acts as the curvature source of the background. To compute the back-reaction on $h_{\mu\nu}$, we use the semiclassical gravity approach and break $h_{\mu\nu}$ into two pieces:
@@
where $h_{\mu\nu}^{\mathrm{(cl)}}$ denotes the classical part, while $h_{\mu\nu}^{\mathrm{(q)}}$ denotes the quantum one with a similar decomposition for $G_{\mu\nu}^{(1)}$. Since $h_{\mu\nu}^{\mathrm{(cl)}}$ is the solution of the linearized Einstein equations sourced by the classical stress-energy tensor, then Eq. (42) implies that the quantum piece of $G_{\mu\nu}^{(1)}$ satisfies
$G_{\mu\nu}^{(1,\,\mathrm{q})}=8\pi\braket{T_{\mu\nu}}.$ (44)
Solving Eq. (44) requires knowledge of the RSET throughout spacetime. However, to capture the asymptotic local contribution of $\braket{T_{\mu\nu}}$ to $h_{\mu\nu}^{\mathrm{(q)}}$, and inspired in the Vaidya metric, we adopt the ansatz
$h_{tt}^{\mathrm{(q)}}=h_{tr}^{\mathrm{(q)}}=-h_{rr}^{\mathrm{(q)}}=\frac{1}{r}q(t-r).$ (45)
The function q(tr)q(t-r) in Eq. (45) is then determined by substituting the ansatz into Eq. (44). After a straightforward evaluation of the linearized Einstein tensor for this ansatz one can show that the leading order terms $\sim 1/r^{2}$ match if
@@
while subleading terms should be fixed by adding subleading contributions to our ansatz. We have added the superscript $\mathrm{(cl)}$} to the curvature scalar $R$ in Eq. (46) to emphasize that it must be computed using the classical background metric.
The complete metric in the asymptotic region $r\to\infty$ outside a spherically-symmetric body of mass $M$ contains a classical part and the quantum correction. We obtain, in spherical coordinates,
$h^{\mu}{}_{\nu}=\eta^{\mu\rho}h_{\rho\nu}=\underbrace{\frac{1}{r}q(t-r)\begin{pmatrix}1&-1&0&0\\ 1&-1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}}_{=h^{\mu}{}_{\nu}^{\mathrm{(q)}}}+\underbrace{\frac{2M}{r}\begin{pmatrix}-1&0&0&0\\ 0&1&0&0\\ 0&0&0&0\\ 0&0&0&0\end{pmatrix}}_{=h^{\mu}{}_{\nu}^{\mathrm{(cl)}}}+\mathcal{O}\left(\frac{1}{r^{2}}\right),$ (47)
@@
At first glance, Eq. (47) might suggest a quantum-corrected gravitational wave. However, several crucial features distinguish it from its classical counterpart. First, the propagating mode in Eq. (47) appears even in spherically symmetric configurations, whereas classical gravitational waves are absent due to Birkhoff’s theorem. Second, the solution does not exhibit the usual transverse polarization. Third, the flux of energy produced by the quantum-corrected metric vanishes at large $r$. Indeed, the gravitational stress-energy pseudo-tensor $t_{\mu\nu}$ is given by [weinberg],
@@
where the superscript $(1)$ denotes the corresponding linearized expression and the superscript $(2)$ the quadratic part in $h_{\mu\nu}=h_{\mu\nu}^{\mathrm{(cl)}}+h_{\mu\nu}^{\mathrm{(q)}}$, a straightforward calculation yields $t^{\mu}{}_{\nu}=\mathcal{O}\left(\frac{1}{r^{3}}\right)\,$. Hence, since $t^{t}{}_{r}$ decays faster than $1/r^{2}$ as $r\to\infty$, no net energy flux is transported to infinity by the quantum-corrected metric.
@@
$\frac{\mathrm{d}^{2}S^{r}}{d\tau^{2}}$ $=\frac{2M}{r^{3}}S^{r}\,$, (49)
$\frac{\mathrm{d}^{2}S^{\theta}}{d\tau^{2}}$ $=\left(-\frac{M}{r^{3}}+\frac{\dot{q}(t-r)}{2r^{2}}\right)S^{\theta}\,$,
$\frac{\mathrm{d}^{2}S^{\varphi}}{d\tau^{2}}$ $=\left(-\frac{M}{r^{3}}+\frac{\dot{q}(t-r)}{2r^{2}}\right)S^{\varphi}\,$,
where, for consistency, we omitted terms proportional to $q(t-r)/r^{3}$. The contributions proportional to the source mass $M$ are the well-known results corresponding to the Schwarzschild geometry, while the terms proportional to $\dot{q}$ in the angular components describe a tidal effect transverse to the propagation direction. Therefore, although the quantum correction to the metric has a polarization tensor that does not represent a genuine gravitational wave, its effect on massive test particles is analog to that of a classical wave.
Finally, we note that, being of order $1/r$, the quantum-corrected metric seems to be of the same order of magnitude than its classical counterpart at large distances. However, restoring conventional units makes the quantum correction proportional to $\ell_{\text{Planck}}^{2}$, so that the classical contribution dominates. An explicit example is discussed in Sec. VI.